# ESIGMAHM: An Eccentric, Spinning inspiral-merger-ringdown waveform model with Higher Modes for the detection and characterization of binary black holes

### v1

@@ Appendix C Model computational cost
We sample $10,000$ binary black hole source configurations and measure the evaluation time of IMRESIGMAHM for each of them, starting from an initial GW frequency of $10$Hz. We choose $1000$ binary configurations for $10$ distinct values of total mass between $10-100M_{\odot}$. We sample mass-ratios uniformly between $[1,6]$, component spins (both independently) between $[-0.9,0.9]$, initial eccentricity between $[0,0.3]$, mean anomaly between $[0,2\pi)$ and inclination angle between $[0,\pi]$. The resulting distributions of evaluation times, along with their median values, for each binary total mass are shown in Fig. 14. We report the evaluation times with IMRESIGMAHM, and also with IMRESIGMA using only the dominant $\ell=m=2$ modes.
Figure 14: Evaluation cost of IMRESIGMAHM and IMRESIGMA generated from an initial GW frequency of $10$Hz is shown as a function of binary total mass. For each value of binary total mass, we sample evaluation times over a uniform distribution of mass-ratio, component spins, eccentricity, mean anomaly and inclination angle. The corresponding median evaluation times are also indicated by the respective markers. The study was performed on an AMD EPYC 7352 processor operating at 2.3 GHz.
For sources with total mass above $30M_{\odot}$ the median evaluation time of IMRESIGMAHM (IMRESIGMA) remains below $1$ $(0.3)$ second, and decreases significantly for higher masses. For the lowest mass of $10M_{\odot}$ the evaluation time can reach up to $\sim 10$ $(\sim 2)$ seconds. It is to be noted that these reported time measurements are with minimal optimizations, and we endeavour to improve upon them in our implementation.
@@ Appendix D NR comparison tables
Table 2: Details of all the quasi-circular NR simulations used in the paper (publicly available in SXS catalog [81]) such as mass ratio $(q)$, reference eccentricity $(e_{\rm{ref}})$, $z$-component of dimensionless spins ($\chi_{1_{z}}$, $\chi_{2_{z}}$), number of orbits $(N_{\rm{orbits}})$, and the corresponding match ($\mathcal{M}$) with IMRESIGMA with dominant $\ell=|m|=2$ modes only.
@@
40 SXS:BBH:0222 1 $<10^{-4}$0.3 $<10^{-6}$ 23.6 98.5
41 SXS:BBH:0046 3 $<10^{-3}$0.5 -0.5 14.4 92.9
42 SXS:BBH:2036 4 $<10^{-3}$ $<10^{-4}$0.4 20.1 94.8
Table 3: Details of all the eccentric NR simulations used in the paper (publicly available in SXS catalog [81]) such as mass ratio $(q)$, reference eccentricity $(e_{\rm{ref}})$, $z$-component of the dimensionless spins ($\chi_{1_{z}}$, $\chi_{2_{z}}$), number of orbits $(N_{\rm{orbits}})$, and the corresponding match ($\mathcal{M}$) with IMRESIGMAHM.

### v2

@@ Appendix C Model computational cost
We sample $10,000$ binary black hole source configurations and measure the evaluation time of IMRESIGMAHM for each of them, starting from an initial GW frequency of $10$Hz at a sampling rate of $4096$Hz. We choose $1000$ binary configurations for $10$ distinct values of total mass between $10-100M_{\odot}$. We sample mass-ratios uniformly between $[1,6]$, component spins (both independently) between $[-0.9,0.9]$, initial eccentricity between $[0,0.3]$, mean anomaly between $[0,2\pi)$ and inclination angle between $[0,\pi]$. The resulting distributions of evaluation times, along with their median values, for each binary total mass are shown in Fig. 14. We report the evaluation times with IMRESIGMAHM, and also with IMRESIGMA using only the dominant $\ell=m=2$ modes.
Figure 14: Evaluation cost of IMRESIGMAHM and IMRESIGMA sampled at $4096$Hz generated from an initial GW frequency of $10$Hz is shown as a function of binary total mass. For each value of binary total mass, we sample evaluation times over a uniform distribution of mass-ratio, component spins, eccentricity, mean anomaly and inclination angle. The corresponding median evaluation times are also indicated by the respective markers. The study was performed on an AMD EPYC 7352 processor operating at 2.3 GHz.
For sources with total mass above $30M_{\odot}$ the median evaluation time of IMRESIGMAHM (IMRESIGMA) remains below $1$ $(0.3)$ second, and decreases significantly for higher masses. For the lowest mass of $10M_{\odot}$ the evaluation time can reach up to $\sim 10$ $(\sim 2)$ seconds. The inspiral mode generation dominates the computational cost, taking roughly twice the time taken for the ODE integration of the inspiral orbital dynamics for IMRESIGMA, and becoming about 10 times that for IMRESIGMAHM. The generation and attachment of the plunge-merger-ringdown piece generated via NRSur7dq4 takes 0.5-7 times the inspiral ODE integration depending on the initial parameters and the mode-content of the waveform, but the inspiral mode generation always dominates the cost. We expect better parallelization of mode generation to further mitigate the overall computational cost of the model. Additionally, we also plan to develop surrogate models for IMRESIGMA to address this point. It is to be noted that these reported time measurements are with minimal optimizations, and we endeavor to improve upon them in our implementation.
@@ Appendix D NR comparison tables
Table 2: Details of all the quasi-circular NR simulations used in the paper (publicly available in SXS catalog [82]) such as mass ratio $(q)$, reference eccentricity $(e_{\rm{ref}})$, $z$-component of dimensionless spins ($\chi_{1_{z}}$, $\chi_{2_{z}}$), number of orbits $(N_{\rm{orbits}})$, and the corresponding match ($\mathcal{M}$) with IMRESIGMA with dominant $\ell=|m|=2$ modes only. The ordering of the NR simulations is done first by increasing mass ratio and then by increasing the value of the dimensionless spin parameter corresponding to the primary object.
@@
41 SXS:BBH:0110 5 $<10^{-3}$ 0.5 $<10^{-6}$ 24.2 99.6
42 SXS:BBH:1463 5 $<10^{-3}$ 0.6 0.2 24.4 99.3
Table 3: Details of all the eccentric NR simulations used in the paper (publicly available in SXS catalog [82]) such as mass ratio $(q)$, reference eccentricity $(e_{\rm{ref}})$, $z$-component of the dimensionless spins ($\chi_{1_{z}}$, $\chi_{2_{z}}$), number of orbits $(N_{\rm{orbits}})$, and the corresponding match ($\mathcal{M}$) with IMRESIGMAHM. The ordering of the NR simulations is done first by increasing mass ratio and then by increasing the value of the dimensionless spin parameter corresponding to the primary object.